Solve the following system by the addition method.
8x+3y=8
8x-5y=8
x=
y=

Answer by mananth(16946)

8x+3y=8
8x-5y=8
multiply second equation by -1
we get -8x+5y=-8
add it to first equation
we get 8x-8x+3y+5y=8-8
8y=0
/8
y=0
plug value of y in any equation
we get
x=1
m.ananth@hotmail.ca